OpenStudy (anonymous):

The vertex of the graph of a quadratic function is (–2, –9) and the y-intercept is (0, –5). Which of the following is the equation of the function? A. y = x2 +10x + 2 B. y = x2 +10x – 2 C. y = x2 – 2x – 5 D. y = x2 + 4x – 5

OpenStudy (anonymous):

All you have to is to simply take one the points given, and substitute for x and for y and see whether they make sense for example, let's take (-2,-9) where -2 = x and -9 = y, now let's see if A. would make sense : y = x^2 + 10x +2 meaning that y = (-2)^2 + 10(-2) +2 so y = 4 -20 +2 = -14 which is incorrect since we were supposed to get y=-9. hence, A. can be ruled out Similarly, use the same method for the rest of the options and you will get the right answer

OpenStudy (anonymous):

Thanks for showing me..